UT 20131215 LUCI2 Nighttime Commissioning

Observer: JHeidt, APramskiy, DThompson
Telescope Operator: GBechetti
LBTO Support: TSargent, DMiller, JHill

Plan for the night

Same as for last night:
- AGw-transform
- LS-acqusition
- MOS-acqusition

Summary

- Suffered from poor seeing

- Took AGw-transform data

- Fiddled with LS-acquistion, still won't work, but know at least whats the problem

- Ghosts/reflection-data

- ZP-measurements

- Twilight flats in J, Br_gam, H2

Preparations

- a bug in the MOS config-file was fixed (persistance mask => persistence mask)
- Alexander fixed the bug in the LS-acq. tool (skysubtracted image had wrong header)
- Michelle provided further scripts

Details

Pls. set FW2 to clear and ini FW2 before new SW version is deployed

Dates are in UT, files are luci2.20131214.iiii

22:40 RON-data, minimum DIT, lir (#2-31), mer (32-61), file #1 crap

23:30 Handover of telescope, Geno prepares for twilight flats

Twilight flats

0062 0071 Br_gam 10x2.75 LIR/NORMAL 11k ADU
0072 0081 H2 10x2.75 LIR/NORMAL 8k ADU
0082 0091 Br_gam 10x2.75 LIR/NORMAL 6k ADU
0092 0101 H2 10x2.75 LIR/NORMAL 5k ADU
0102 0111 Br_gam 10x2.75 LIR/NORMAL 2k ADU
0112 0121 H2 10x2.75 LIR/NORMAL 2k ADU

Two faults on FW2, LUCI2 apparently can't make or send TRACK mode presets!
So, we missed K flats.

0122 0131 J 10x2.75 LIR/NORMAL 9k ADU
0132 0141 J 1x2.75 Dithered 8...4k
0142 0151 J 1x2.75 Dithered 4...3k
0152 0162 J 1x2.75 Dithered ~2.5k (one extra)

00:20 Sunset, weather is great, clear, low wind, low humidity (30%), T_ambient = -3.5deg, T_mirror = -2.9deg

01:00 Geno prepares for AGw-transform data taking on both sides, seeing is poor 1.5-2.5" on guider

AGw-transform determination

01:45 DMiller startet some AGw spin-tests

Configure for AGw transform calibration
- pointing mask
- blind+clear until ready

0163 K 3x4.0s LIR/INT test exposure for transform check
0164 K 4x3.0s --"--

0165 0169 K 4x3.0s with blind filter as "dark" to see circle of dots

transform_collect on Stone_A1_15
0172 "sky image"
0179 script failed, couldn't find info file?
0180 gs=8 restarted transform_collect with START=5
0181 gs=2
0182 gs=4
script failed 3x, couldn't find info file? spot too dim? gs=4 (START=7)
0183 gs=5 (START=8)
0184 gs=4 failed again?
0185 gs=2 (START=10)
0186 gs=2 repeated exposure?
0187 gs=8
0188 gs=9
0189 gs=10
0190 gs=7
0191 gs=3 PA=-119
0192 gs=3 PA=-97
0193 gs=7 (should be about 15 measurements here)
0194 gs=10
0195 gs=9
0196 gs=8
0197 gs=2
0xxx gs=4 GEIRS error 63 and spots too dim again!
0198-0215 recovery exposures
0216 gs=4

Change authorization to just LUCI2

03:07 JHill joined by video

LS-acquisition

05:23 Preset to BS9108, check pointing, than active preset and collimate
seeing still bad ~2"

05:40 Test on whether alignment of one star at PA=227 works now
#217-220 crap, #221 source => #222 verification image,
Source = 1108,1191, move-to = 977,896, moved to 1269, 876 (#222) => crap

06:00 Preset BS9108, PA=0, start from scratch, source (#223), verification (#224)
source=1057,1020, requested=969,912, found=974,907 looks OK +-5pix
With sky subtraction: #225 sky, #226 source, source=976,825, requested=895,908,
verification #227, found=893,904 => seems Ok too

06:18 Preset BS9108, PA=270, GS=9, source #228, verification #229,
source=1051,1008, requested=971,910, found=951,1082
Shift in x (delta) seems OK, shift in y (alpha) has wrong sign

06:37 Preset to BS9108, PA=180, GS=9, #230, source=1050,1052,
requested=978,915, verification #231, found=1115,1215, wrong sign in alpha,delta

06:52 Preset to BS9108, PA=90, GS=4, #232, source=1047,1019
requested=965,904, verification #233, found=1166,938, wrong sign in delta

Change of signs, alpha,delta
PA=0: OK, OK
PA=90: OK; wrong
PA=180:wrong,wrong
PA=270:wrong,OK

Unclear whats wrong, need to talk to Ingo.
=> In fact, all is done in d(alpha)/d(delta), ie can not work
==> must be in detxy-system => AI IStilz has to fix that rapido

Focuscheck N1.8-camera

08:40 Check focus N1.8-camera, Focus=-285, images in KKsHJ, also be useful as ZPs
K, #251, fwhm=5.6,6.3pix
Ks, #252, fwhm 4.5,5.3pix
H, #253, fwhm=6.4,6.4pix
J, #254, fwhm=5.5,5.5pix
Set N1.8-camera, focus=-285
K, #255, fwhm=5.7,5.5pix
Ks, #256, fwhm=5.2,5.1pix
H, #257, fwhm=5.7,5.4pix
J, #258, fwhm=7.3,6.3pix
=> no obvious difference

Photometric ZPs

09:08 Its photometric outside, ZP's in JHKKs/N3.75 at BS9126, AM=1.004, #259-278

10:03 ZP's in JHKKs/N3.75 at BS9108, AM=1.9, #279-98

11:25 Execute JKurk script M242SDSS1110.xml to take "practice" MOS
alignment data. Failed (same reason as below)

11:33 Error in the supplied coordinates somehow?
Sun Dec 15 11:28:19.230 2013 Preset alert for right LUCIFER mode ACTIVE: AGw command status:
AGW2: Requested position is Out of Range, Sun Dec 15 11:28:19.279 2013 Preset alert for right LUCIFER mode ACTIVE:
error: setProbePosition: couldn't move probe to position (SFP) (x,y,z): -1.43039e+08, 1.10106e+08, -1.57288e+13 mm.

Ghosts

11:43 M5_1311 @ 12:00:52 +06:36:51 R=4.6mag
gs=1 is at radius 314", some dithers for this will likely take it
out of the patrol field.

0292 3x4 .0s K
0293 3x20.0s K
0294 3x20.0s K

0295 3x20.0s H
0296 3x20.0s H

0297 3x20.0s J
0298 3x20.0s J

0299 3x20.0s z
0300 3x20.0s z

0301 3x20.0s blind in FW1

0302 3x20.0s P_gam
0303 3x20.0s P_gam

0304 3x20.0s FeII
0305 3x20.0s FeII

0306 3x20.0s H2
0307 3x20.0s H2

Test of offset-only script - it sent a preset!

blind+blind filters
0308 0317 10x10.0s LIR/NORMAL to see persistence from the v.bright star
0318 1 x100.0s MER/NORMAL

12:40 Test of offset-only script - it sent a preset!

12:40:26 Error while trying to move a mask from the focal plane to turnout.
However, when this happened there was no mask in the FPU.

0319 1 x100.0s MER/NORMAL (oops, forgot to move in blind filters?)

Photometric ZPs

12:49 Photometric standard BS9126
Error while commiting to readout manager
Abort button on READ2 GUI just gives "Empty GEIRS response" popup

320 readout I could not abort

UT 13:01 restart BS9126 script
again same error. Config for 1x3.0s and manual read...error 63, repeats.

UT 13:07 restart BS9126 script
0321 0336 5x10.0s various filters

0337 0356 test of Jochen's offset-only script

Twilight flats

UT 13:35 Move to BS9149 for twilight flats
0357 test 1x2.75 H ~3500 counts
0358
0359 K test
0360 0363 more tests/recovery of FW2
Lost BB morning twilight flats to FW2 problems

Moving to zenith away from the sun for NB flats.
Too bright for NBs except BrG and H2 (already done once)

0364 0373 10x2.75 LIR/NORM Br_gam 3.5k
0374 0383 10x2.75 LIR/NORM H2 4k

0384 0393 10x2.75 LIR/NORM Br_gam 6.5k
0394 0403 10x2.75 LIR/NORM H2 7.5k

0404 0413 10x2.75 LIR/NORM Br_gam 8.7k
0414 0423 10x2.75 LIR/NORM H2 9.5k

14:30 Closing

-- JochenHeidt - 15 Dec 2013
Topic revision: r10 - 16 Dec 2013, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback